# Optimal Projection for 3D Gaussian Splatting

### v2

@@ 6 Experiments
Table 1: Quantitative evaluation of our method compared to previous work, computed over three datasets. The Mip-NeRF360 [2] dataset encompasses both Indoor and Outdoor scenarios, while Tanks. and Deep. indicates the average results over the Tank&\&&Templates [17] and Deep Blending [12] datasets, respectively. The term Average denotes the mean value of measurements across all scenarios in these datasets.
@@
Figure 5: We show comparisons of ours to previous methods and the corresponding ground truth images from held-out test views. Due to space constraints, we have selected methods such as 3D-GS [16] and Mip-NeRF360 [2], which are comparable in terms of metrics, for qualitative comparisons. The scenes are, from the top down: Truck and Train from Tanks&Temples [17]; Playroom, DrJohnson from the Deep Blending dataset [12] and Room, Bonsai, Counter and Stump from Mip-NeRF360 dataset [2]; Non-obvious Differences in quality highlighted by arrows/insets.
To further validate the effectiveness of the Optimal Gaussian Splatting derived from theoretical analysis, we conducted a series of experiments, comparing it with the original 3D-GS [16] and some current state-of-the-art methods.
6.1 Implementation
6.2 Datasets
We tested our algorithm on a total of 13 real scenes which are the same as those used in the original 3D-GS [16] paper. In particular, we evaluated our approach on the complete set of scenes featured in Mip-NeRf360 [2], which currently represents the state-of-the-art in NeRF rendering quality. Additionally, we tested our method on two scenes from the Tanks & Temples dataset [17] and two scenes provided by Hedman et al.  [12]. The selected scenes exhibit diverse capture styles, encompassing both confined indoor environments and expansive, unbounded outdoor settings. In other words, our evaluation includes a wide range of scenes to assess the performance of our method comprehensively.
6.3 Comparisons
We compare our algorithm with 3D-GS [16] and three state-of-the-art NeRF-based approaches: Mip-NeRF360 [2], InstantNGP [26] and Plenoxels [7]. All methods were configured using the same settings as outlined in the 3D-GS paper to control variables, preventing the introduction of variables other than our Optimal Splatting method.
6.4 Results
@@
We adopt a train/test split for datasets following the methodology proposed by Mip-NeRF360 [2]. Specifically, every 8th photo is reserved for testing, ensuring a consistent and meaningful basis for comparisons to generate error metrics. Standard metrics such as PSNR, LPIPS [43], and SSIM, widely used in the literature, are employed for evaluation; refer to Table 1 for details. It can be observed that, while the 3D-GS method significantly outperforms MipNeRF360 in rendering speed, its rendering quality slightly falls behind MipNeRF360, with a lower average PSNR across the 13 scenes. This might be attributed to MipNeRF360 not employing significant approximations throughout its entire pipeline. However, by replacing the original projection method of 3D-GS with our Optimal Projection, which incurs smaller errors, we not only maintain rendering speed far surpassing MipNeRF360 but also achieve a quality improvement. Furthermore, in terms of average results across the 13 scenes, our approach outperforms others, including 3D-GS and several NeRF-based methods, in PSNR, SSIM, and LPIPS.
@@
Due to space constraints, for our qualitative experiments, we selected only two methods, 3D-GS and MipNeRF360, from Table 1, which have comparable metrics, as our baselines.
Please see Figure 5. From the comparative images, it is evident that our algorithm, compared to 3D-GS, reduces artifacts and defects while preserving more details. In the Truck scene, we address the absence of the steering wheel in the reconstruction by 3D-GS (indicated by the red arrow), while avoiding the pillar fragmentation introduced by MipNeRF360 (indicated by the blue arrow). Additionally, in the Train scene, we maintain the geometry of the train’s iron railing while avoiding the black floaters present in 3D-GS. In the PlayRoom scene, we achieve relatively better wire geometry. In the DrJohnson scene, our algorithm neither exhibits the fragmented appearance seen in 3D-GS nor the blurred details observed in MipNeRF360. In the Room and Bonsai scenes, we respectively have better stems in the flowerpot and improved details on the small white objects on the floor compared to 3D-GS. In the Counter scene, our floor texture is optimal. In the Stump scene, our tree branches do not fracture as observed in 3D-GS.
7 Discussion and Conclusions
Because Gaussian functions remain Gaussian after affine transformations, convolutions, and similar operations, but this is not necessarily the case for projection transformations, the 3D Gaussian Splatting process employs a Taylor first-order expansion to approximate this projection transformation as an affine transformation. However, such approximation inevitably introduces errors, which can result in artifacts, distortions, and inaccuracies in the learning comprehension of the scene as well as the generation of novel view images. This paper begins by examining the Taylor expansion remainder of the local affine approximation, obtaining the mathematical expectation of the error. This expectation is a function of the mean of the Gaussian function. By finding critical points, it is determined that the function attains its minimum when the projection of the mean of the Gaussian function on the plane coincides with the projection of the camera center on the plane. The visualization of this function is depicted in Figure 2. Based on the conclusions drawn from this theoretical analysis, we propose a more optimal projection method with reduced error, referred to as Optimal Gaussian Splatting. Subsequently, extensive experiments are conducted to further validate the superiority of this projection approach.
The original 3D-GS [16], despite exhibiting significantly superior performance compared to the NeRF-based MipNeRF360 [2] method, may suffer a slight compromise in quality due to errors introduced by projection, as opposed to MipNeRF360 [2]. In contrast, our Optimal Gaussian Splatting, by minimizing projection errors, unlocks a portion of the untapped potential of 3D-GS. This demonstrates that the representation of scenes using 3D-GS can achieve state-of-the-art (SOTA) levels in both quality and performance. Additionally, it is noteworthy that our entire error analysis process only considers the influence of Gaussian mean on the error function while treating the covariance as a constant. Hence, future research could explore optimizing 3D-GS further by considering the Gaussian covariance in the analysis.
Furthermore, our method, throughout its entire process, is only dependent on the transformation from image pixels to the cutting plane (corresponding to Equation 23) and its relationship with the image representation. Unlike the original 3D-GS, where the $z=1$ plane tends to favor perspective images due to the perspective projection viewport transformation involving the NDC’s $z=1$ plane, our approach offers more versatility. This makes our method potentially applicable to various image representations. However, this versatility may come with a slight performance trade-off or require appropriate code implementation to avoid extending training time due to the additional operation of pixel-to-tangent-plane transformation, as shown in Equation 23.
In summary, our work represents the first theoretical analysis that focuses on the projection error of 3D Gaussian Splatting. We have introduced a novel Splatting method Optimal Gaussian Splatting aimed at reducing projection errors, and through both theoretical analysis and empirical validation, demonstrated the superiority of this approach, as shown in Figure 1 and Figure 5.

### v4

@@ 6 Experiments
To validate the effectiveness of the Optimal Gaussian Splatting derived from theoretical error analysis, we conduct a series of experiments, comparing it with the original 3D-GS [17] and some current state-of-the-art methods.
6.1 Implementation
6.2 Datasets
We test our algorithm on a total of 13 real scenes which are the same as those used in the original 3D-GS [17]. In particular, we evaluate our approach on the complete set of scenes featured in Mip-NeRF360 [2], which currently represents the state-of-the-art in NeRF rendering quality. Additionally, we test our method on two scenes from the Tanks & Temples dataset [18] and two scenes provided by Hedman et al.  [14]. The selected scenes exhibit diverse capture styles, encompassing both confined indoor environments and expansive, unbounded outdoor settings.
6.3 Results
Table 1: Quantitative evaluation of our method compared to previous work across three datasets. The Mip-NeRF360 [2] dataset includes both indoor and outdoor scenarios. Tanks. and Deep. indicates the average results over the Tanks&\&&Templates [18] and Deep Blending [14] datasets. Average represents the mean value across all scenarios in these datasets.
@@ 6.3 Results
Figure 4: We show comparisons of our method to previous methods and the corresponding ground truth images from held-out test views. The scenes are, from the top down: Truck from Tanks&\&&Temples [18]; Playroom from the Deep Blending dataset [14] and Bonsai, Counter from Mip-NeRF360 dataset [2]. Differences in quality highlighted by arrows/insets.
Figure 5: We derive the projection error function with respect to image coordinates $(u,v)$ and focal length through the transformation between image and polar coordinates $(\theta_{\mu},\phi_{\mu})$. We visualize this function, showing 3D-GS rendered images with long and short focal lengths, followed by the corresponding error functions.
We compare our algorithm with 3D-GS [17] and three state-of-the-art NeRF-based approaches: Mip-NeRF360 [2], InstantNGP [28] and Plenoxels [9]. All methods are configured using the same settings as outlined in the 3D-GS [17] to control variables, preventing the introduction of variables other than our Optimal Splatting method.
@@
We adopt a train/test split for datasets following the methodology proposed by Mip-NeRF360 [2]. Specifically, every 8th photo is reserved for testing, ensuring a consistent and meaningful basis for comparisons to generate error metrics. Standard metrics such as PSNR, LPIPS [45], and SSIM are employed for evaluation. We report quantiative results in Table 1. While the 3D-GS significantly outperforms MipNeRF360 in rendering performance, its rendering quality slightly falls behind MipNeRF360, with a lower average PSNR across the 13 scenes. This might be attributed to MipNeRF360 not employing significant approximations throughout its entire pipeline. However, by replacing the original projection method of 3D-GS with our optimal projection, which incurs smaller errors, we not only maintain rendering performance far surpassing MipNeRF360 but also achieve a quality improvement.
@@
For our qualitative experiments, we select only two methods, 3D-GS and MipNeRF360, as our baselines. It can be observed that our method is capable of generating more realistic details, with fewer artifacts compared to 3D-GS [17]. And this is precisely the superiority brought about by our projection method, which results in smaller errors. Furthermore, We see that in some cases Mip-NeRF360 has remaining artifacts that our method avoids (e.g., black lines - in Playroom).
6.4 Discussions
Impacts of Decreasing Focal Length
Based on Equation 9, we further derive the projection error function with respect to image coordinates $(u,v)$ and focal length $f_{x},f_{y}$ through the coordinate transformation between image coordinates and polar coordinates $(\theta_{\mu},\phi_{\mu})$. As illustrated in Figure 5, the peak of projection error mainly occurs at the image edges under a long focal length, resulting in artifacts at the image edges. As the focal length decreases, the field of view expands, leading to more Gaussians deviating from the projection center and an overall increase in projection error. In such cases, 3D-GS exhibit more artifacts, such as needle-like structures or cloud-like large Gaussians, which obscure parts that perform well under long focal length, thereby significantly degrading the overall image quality. Since our method employs a central radial projection for the projection plane, theoretically, we should not encounter such defects. To validate this perspective, we conduct experiments at various focal length reduction ratios (reduction ratios of ×$\times 0.2$ and ×$\times 0.3$).
Table 2: We quantitatively compare our method with the original 3D-GS [17] using focal lengths of 0.2 and 0.3 times the original. Six scenes were selected from three datasets: Mip-NeRF360 [2], Tank &\&& Templates [18], and Deep Blending [14]. The scenes include Train and Truck from Tanks &\&& Templates, DrJohnson and Playroom from Deep Blending, and Room and Treehill from Mip-NeRF360.
Metric Focal Method Train Truck Dr Johnson Playroom Room Treehill Average
[…]
Ours 16.60 16.70 23.54 23.29 23.67 18.98 20.46
$\times$0.3 3D-GS [17] 12.85 14.89 23.36 22.51 25.00 18.35 19.49
Ours 17.52 18.07 24.89 26.40 25.87 19.87 22.10
[…]
Ours 0.565 0.537 0.701 0.752 0.777 0.435 0.628
$\times$0.3 3D-GS [17] 0.471 0.529 0.740 0.791 0.812 0.437 0.630
Ours 0.588 0.605 0.762 0.844 0.834 0.472 0.684
[…]
Ours 0.277 0.250 0.216 0.203 0.178 0.385 0.251
$\times$0.3 3D-GS [17] 0.331 0.276 0.207 0.213 0.154 0.397 0.263
Ours 0.243 0.217 0.201 0.182 0.158 0.385 0.231
Figure 6: We compare our method to the original 3D-GS [17] with different focal length reductions. Two focal lengths, 0.2 and 0.3 times the original, are tested. Ground truth ×$\times 1$ in the left images indicates no focal length reduction. The red-boxed areas in the right images correspond to the same regions in the left images.
Figure 7: Renderings of three scenes using our projection under three camera models. From left to right: the ground truth of the original perspective, wide-angle rendering with a pinhole camera, fisheye camera rendering, and the panorama.
For quantitative experiments, considering that there is no ground truth for large field-of-view in the 13 scenes from the 3D-GS [17], we apply a focal length mask. This mask selects central parts of the rendered images with corresponding ground truth based on the focal length scaling factor for metric calculation. The results are presented in the Table 2. Whether it is at ×$\times 0.2$ or ×$\times 0.3$ reduction ratios, our method outperforms the original 3D-GS significantly across various scenes in terms of all metrics. As the focal length decreases from ×$\times 0.3$ to ×$\times 0.2$, 3D-GS [17] experiences a significant degradation due to increased projection errors, while our method demonstrates greater robustness.
For quantitative experiments, we randomly select three scenes, each with two different focal lengths for demonstration. Additional experimental results can be found in the supplementary materials. It is evident from the Figure 6 that our method exhibits greater robustness across various focal length settings, mitigating the artifacts introduced by projection in the original 3D-GS.
Adaptability to Various Camera Models
By straightforwardly modifying the transformation from image space to camera space, we have achieved adaptation to various camera models, with the caveat that the original 3D-GS’s projection struggles to support these camera models. Figure 7 illustrates that our projection, in addition to conforming to the original pinhole camera according to Equation 16, can also accommodate fisheye cameras and generate panoramas.
Limitation
Due to our method’s independence from the image plane, it exhibits adaptability to various camera models. However, the slight increase in training time is a consequence of the subsequent transformation to the image plane after projection onto the tangent plane, rather than a direct projection onto the image plane. Fortunately, this method avoids the need for dense points sampling, ensuring the algorithm’s time complexity remains unchanged and rendering performance stays unaffected. Moreover, it may be possible to mitigate the impact on training time, particularly concerning the corresponding backward propagation, through the implementation of optimized CUDA code.
7 Conclusion
Through theoretical analysis, we establish the correlation between the projection error and the Gaussian mean in 3D-GS. By optimizing the error function, we discover that the projection error is minimized when projecting on the plane perpendicular to the line connecting the Gaussian mean and the camera center. Leveraging this insight, we propose a novel projection that yields high-quality images without compromising rendering performance and can be easily adapted to various camera models through simple modifications.
In this work, our primary focus is on analyzing the relationship between the projection error and the mean of the 3D Gaussian under the assumption of a constant covariance. The influence of Gaussian covariance on projection requires further discussion in future work. It may be possible to further explore the potential of Gaussian Splatting as an explicit representation technique for scenes.
Appendix 0.A Details of Error Analysis
Unit Vector Simplification
We have proved that the composition of the transformation $\varpi$, projecting onto the unit sphere, and the transformation $\varphi$, projecting onto the tangent plane, is equivalent to the single transformation $\varphi$ directly projecting onto the tangent plane:
$\begin{split}\left(\varphi\circ\varpi\right)\left({\mathbf{x}^{{}^{\prime}}}\right)&=\varphi\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\\ &={\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\left(\mathbf{x_{0}}^{\top}\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\right)^{-1}\\ &={\mathbf{x}^{{}^{\prime}}}\left(\mathbf{x_{0}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1}\\ &=\varphi\left({\mathbf{x}^{{}^{\prime}}}\right)\text{.}\end{split}$ (19)
Therefore, we can simplify the derivation of the error function $\epsilon$ by projecting all relevant points onto the unit sphere as unit vectors (Equation 8 in the main paper).
Closed-Form Expression for the Error Function
This section provides a comprehensive derivation of the integral expression for the error function as presented in the main paper. Upon expressing $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ as unit vectors and substituting them into Equation 6 in the main paper, the Jacobian matrix is derived:
$\mathbf{J}=\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\left[\begin{matrix}\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&0&-\frac{\sin{\left(\phi_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}\\ 0&\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&\frac{\sin{\left(\theta_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}\\ 0&0&0\end{matrix}\right]\text{.}$ (20)
Substituting this Jacobian matrix into Equation 5 in the main paper yields the derivation of the Taylor expansion remainder term:
And substituting this expression into Equation 7 in the main paper results in the integral expression (Equation 9 in the main paper). The main paper provides the integral expression (Equation 9 in the main paper) and the function graph (Figure 2 in the main paper) for the error function, without presenting the closed-form expression after integration. We provide the closed-form expression as follows:
$\begin{split}\epsilon&\left(\theta_{\mu},\phi_{\mu}\right)=\frac{1}{{16\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}}}\\ &\Big{\{}8\left(\left(2\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}-\pi\right)\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+2\right)\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan^{2}{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+\\ &\left.\left(-4\cdot\left(2\cos{\left(2\theta_{\mu}\right)}+\pi\right)\sin^{2}{\left(\phi_{\mu}\right)}+\pi\left(2\cos{\left(2\theta_{\mu}\right)}+\pi\right)+4\cos{\left(2\theta_{\mu}\right)}+2\pi\right)\right.\\ &\left.\sin^{2}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+8\log{\frac{{\left(\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}}{{\left(1-\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\left(1-\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}}}\right.\\ &\left.\log{\frac{{\left(1-\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\left(1-\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\exp{\left(2\sqrt{2}\cos{\left(\theta_{\mu}\right)}\right)}}}{{\left(\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\tan^{2\sin{\left(\theta_{\mu}\right)}}{\left(\theta_{\mu}+\frac{\pi}{4}\right)}\right)}}}\right.\\ &\left.\cos^{3}{\left(\phi_{\mu}\right)}\cos^{3}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.4\pi\left(-4\log{\left(\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}\tan{\left(\phi_{\mu}\right)}+\pi\tan^{2}{\left(\phi_{\mu}\right)}+2\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}-\pi\right)\right.\\ &\left.\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.\left(2\pi\left(\left(4+2\pi+16\sqrt{2}\right)\sin^{2}{\left(\theta_{\mu}\right)}-2+\pi\right)\cos^{2}{\left(\phi_{\mu}\right)}+\left(2\pi-4\right)\cos{\left(2\theta_{\mu}\right)}-2\pi+\pi^{2}\right)\right.\\ &\left.\cos^{2}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.16\sqrt{2}\log{\frac{{\left(1-\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\left(1-\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\exp{\left(2\sqrt{2}\cos{\left(\theta_{\mu}\right)}\right)}}}{{\left(\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}}}\right.\\ &\left.\cos^{3}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.64\sin^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}-\right.\\ &\left.16\sqrt{2}\cdot\left(1+\sqrt{2}\right)\sin{\left(\theta_{\mu}\right)}\sin{\left(2\theta_{\mu}\right)}\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &8\pi\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\frac{8\left(-2\pi\cos^{2}{\left(\theta_{\mu}\right)}+\pi+4\right)\cos^{4}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}{\left(\tan{\left(\theta_{\mu}\right)}-1\right)^{2}}\Big{\}}\text{.}\end{split}$ (22)
Appendix 0.B Details of Adaptation for Various Camera Models
In the main paper, it is highlighted that our projection, independent of perspective image plane, allows for the adaptation to various camera models by modifying the transformation from image space to camera space in the rasterization based on the unit sphere (Equation 16 in the main paper) . This section provides a detailed explanation of our method’s adaptation to various camera models.
Fisheye
The design models for fisheye cameras can generally be categorized into four types: equidistant projection model, equisolid angle projection model, orthographic projection model, and stereographic projection model. We explane our projection’s adaptation to fisheye camera models, taking the equidistant projection model as an example.
For a pixel $\left(u,v\right)$ on the image, we cast a ray. According to the transformation between the image space and camera space for the equidistant projection model, we obtain:
$\mathbf{x}_{\text{2D}}=\varphi_{\text{p}}\left(\left[\begin{matrix}\frac{\left(u-c_{x}\right)\sin{\left(\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}\right)}}{f_{x}\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}}\\ \frac{\left(v-c_{y}\right)\sin{\left(\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}\right)}}{f_{y}\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}}\\ \cos{\left(\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}\right)}\end{matrix}\right]\right)$ (23)
where $c_{x}$, $c_{y}$, $f_{x}$, $f_{y}$ denote the intrinsic parameters of the camera model.
Panorama
For a pixel $\left(u,v\right)$ on the image, we cast a ray. According to the transformation between the image space and camera space for panoramic images, we obtain:
$\mathbf{x}_{\text{2D}}=\varphi_{\text{p}}\left(\left[\begin{matrix}\sin{\left(\frac{\pi\left(-W+2u\right)}{W}\right)}\cos{\left(\frac{\pi\left(-\frac{H}{2}+v\right)}{H}\right)}\\ \sin{\left(\frac{\pi\left(-\frac{H}{2}+v\right)}{H}\right)}\\ \cos{\left(\frac{\pi\left(-\frac{H}{2}+v\right)}{H}\right)}\cos{\left(\frac{\pi\left(-W+2u\right)}{W}\right)}\end{matrix}\right]\right)$ (24)
where $H,W$ represent the height and width of the image, respectively.
Then, similar to Equation 17 in the main paper, the 2D Gaussian function values can be obtained for alpha blending to generate the image. The alpha blending process after obtaining the 2D Gaussian is the same as in the original 3D-GS [17].
Figure 8: We show comparisons of our method to the original 3D-GS [17] under various camera models and different focal lengths.
Appendix 0.C Details of Experiments
Additional Results
Table 3-4 list PSNR for our evaluation over all considered techniques and real-world scenes, corresponding to Table 1 in the main paper. Figure 9 illustrates additional qualitative comparisons between our method and other approaches.
Figure 9: We show comparisons of our method to previous methods and the corresponding ground truth images from held-out test views. The scenes are, from the top down: Train from Tanks&Temples [18]; Drjohnson from the Deep Blending dataset [14] and Room, Stump from Mip-NeRF360 dataset [2]. Differences in quality highlighted by arrows/insets.
Impacts of Decreasing Focal Length
Figure 10: Illustration of details of metrics caculation in the quantitative comparisons on the impact of decreasing focal length.
In the quantitative comparison experiments about the impact of reducing focal length, we utilized a focal length mask methodology due to the absence of ground truth for wide-angle images. The detailed procedures are illustrated in Figure 10. The focal mask selects central parts of the rendering with corresponding ground truth based on the focal length scaling factor for metric calculation. For rendering, we crop the corresponding part with available ground truth, then resize the ground truth to match the size of this patch. Finally, we compute the metrics for these two images.
For qualitative experiments, we show additional results in Figure 8. In addition to the original focal length reduction factors of 0.3 and 0.2, we introduce 0.1 times the focal length. Additionally, we generate images with corresponding perspectives using other camera models, while 3D-GS [17] is not presented as it lacks support for these camera models.
Comparison with the Other 3D-GS Methods
We choose Mip-Splatting [44] and Scaffold-GS [22] to compare in the same setting as in Figure 6 of the main paper. For Mip-Splatting [44] and Scaffold-GS [22] still employ conventional projection of 3D-GS, the approximation error remains larger compared to ours, as shown in Figure 11. Additionally, our method can be be seamlessly integrated with theirs to reduce the projection error of these methods.
Figure 11: We show comparisons of our method to the original 3D-GS [17], Mip-Splatting [44] and Scaffold-GS [22] with a large FOV camera. In the figure, Ours(+X) indicates the combination of our method with the X method.
Adaptability to Various Camera Models
Since our method is based on differentiable rasterizer, it can also be used to train on the non-pinhole camera dataset. We illustrate the results obtained from training on the non-pinhole camera dataset Matterport [5] directly, where 3D-GS fails entirely to reconstruct the scene, as shown in Figure 12.
Figure 12: We show the results of our method trained on the non-pinhole camera dataset Matterport [5].
Table 3: PSNR scores for Mip-NeRF360 scenes.
MipNeRF360 Indoor MipNeRF360 Outdoor Avg.
Room Counter Kitchen Bonsai Avg. Bicycle Flowers Garden Stump Treehill Avg.
Plenoxels [9] 27.59 23.62 23.42 24.67 24.83 21.91 20.10 23.49 20.66 22.25 21.68 23.08
INGP-Base [28] 29.27 26.44 28.55 30.34 28.65 22.19 20.35 24.60 23.63 22.36 22.63 25.30
INGP-Big [28] 29.69 26.69 29.48 30.69 29.14 22.17 20.65 25.07 23.47 22.37 22.75 25.59
M-NeRF360 [2] 31.63 29.55 32.23 33.46 31.72 24.37 21.73 26.98 26.40 22.87 24.47 27.69
3D-GS [17] 30.63 28.70 30.32 31.98 30.41 25.25 21.52 27.41 26.55 22.49 24.64 27.21
Ours 31.58 29.05 31.23 32.31 31.04 25.07 21.54 27.15 26.57 22.77 24.62 27.48
Table 4: PSNR scores for Tanks&Temples and Deep Blending scenes.
Tanks&Temples Deep Blending Avg.
Truck Train Avg. Dr Johnson Playroom Avg.
Plenoxels [9] 23.22 18.93 21.07 23.14 22.98 23.06 22.07
INGP-Base [28] 23.26 20.17 21.72 27.75 19.48 23.62 22.67
INGP-Big [28] 23.38 20.46 21.92 28.26 21.67 24.96 23.44
M-NeRF360 [2] 24.91 19.52 22.22 29.14 29.66 29.40 25.81
3D-GS [17] 25.19 21.10 23.14 28.77 30.04 29.41 26.27
Ours 25.19 21.67 23.43 29.14 29.89 29.51 26.47